foxyisthebest9000:

The dimensions of a rectangular prism are shown below: Length: 1 and 1 over 3 feet Width: 1 foot Height: 2 and 1 over 3 feet The lengths of the sides of a small cube are 1 over 3 foot each. Part A: How many small cubes can be packed in the rectangular prism? Show your work. (5 points) Part B: Use the answer obtained in part A to find the volume of the rectangular prism in terms of the small cube and a unit cube. (5 points)

jhonyy9:

please post it in the mathematics subject

supie:

Ok so in order to find part A, ` How many small cubes can be packed in the rectangular prism` we would have to divide the mass of the cubes by the rectangular prism that would tell us how many of `the value of a cube` could go into `the rectangular prism` In other words how much times can \(\frac13ft\) go into the rectangular prism ( when Length: 1 and 1 over 3 feet~Width: 1 foot~Height: 2 and 1 over 3 feet) (\(\frac13ft\) is how big each small cube is [since it's a cube every side has the same length]) So basically it's \(\frac13ft/*rectangular\ prism*\) or \(\frac13fr÷*rectangular\ prism*\)

supie:

So we have the dimensions: \(L=1\frac13ft\\ W=1ft\\ H=2\frac13\)


We know that the `small cube` has \(\frac13ft\) on `each` side. So we use the volume formula which is \(V=L*W*H\) so we plug in the length width and height values \(V = \frac43*1*\frac73 = \frac{28}{9} ft^3\)

supie:

So then we find the small cube's volume \((\frac13)^3 = \frac{1}{27} ft^3\) So then we basically combine them ~divide \(\frac{\frac{28}{9}}{ \frac{1}{27}} = \frac{28}{9}*27 = 28*3\) Which is basically \(\frac{28}{9} ÷ \frac{1}{27} = \frac{28}{9}*27 = 28*3\) Which =? that would be the number of small cubes that could fit inside the rectangular prism that has a length of \(1 \frac13ft\) a width of 1 foot, and height of \(2 \frac13\) which would be your final answer